# Construction of a Basis of Kähler differentials for generic plane branches

### v1

@@ 4.1. Setting for the case $E\neq\emptyset$
Notice that the only modification is to require that $E$ is invariant, in $E$-(i) and $E$-(ii).
@@
A family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}\subset\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $l_{0}\leq l\leq g$, is $E$-terminal if
  • $E$-(a)
@@
Given an $E$-terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ of level $0\leq l<g$, we define an $E$-initial family ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2\nu_{l+1}}^{0})$ of stage 00 (in level $l+1$) by means of equation (5), exactly as in the case $E=\emptyset$.
@@ 4.2. The method for the case $E\neq\emptyset$
When $E=\emptyset$, the terminal family of level 00 was ${\mathcal{T}}_{0}=\{dx,df_{1}\}$. From this family, applying the results in subsections 3.5.2 and 3.5.3 we ended up with the terminal family ${\mathcal{T}}_{g}$.
@@
Thus, we only need to find an $E$-terminal family of some level $l_{0}$. Actually, it suffices to find an $E$-good family ${\mathcal{G}}_{l_{0}}^{1}$ (with $E$-(I) being empty) since in such a case the results in subsection 3.5.3 provide an $E$-terminal family of level $l_{0}$.
@@ 4.2.1. Case $E=\{x=0\}$
which satisfy $\nu(\Omega_{1})=0$, $\nu_{\mathcal{C}}(\Omega_{1})=n$, $\nu(\Omega_{2})=1$, $\nu_{\mathcal{C}}(\Omega_{2})=\beta_{1}+n$. The family $\mathcal{T}_{0}=\{\Omega_{1},\Omega_{2}\}$ is clearly $E$-terminal of level $l_{0}=0$ and thus, we obtain an $E$-nice basis $\mathbf{\Omega}$ by Remark 4.1.
@@ 4.2.2. Case $\delta_{y}=1$ and $\beta_{0}\in(n)$
Notice that $\beta_{0}\in(n)$ implies $\beta_{0}<\beta_{1}$. We have $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$ and hence $\nu_{C}(\overline{\Omega}_{1})\in(n)$, $\nu_{C}(\overline{\Omega}_{2})\in(n)$. We set $l_{0}=0$, $\Omega_{1}^{0}=\overline{\Omega}_{1}$ and $\Omega_{2}^{0}=\overline{\Omega}_{2}$. Applying Lemma 3.5, with $\beta_{l+1}=\beta_{0}$, we obtain $\Omega_{1}^{1}=\Omega_{1}^{0}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{0}),<}^{0}$ where $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{0})$. The family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is an $E$-good family of level 00. Moreover, it satisfies $\nu(\Omega_{1}^{1})=\delta_{x}$ and $\nu(\Omega_{2}^{1})=1$. Remark 4.1 guarantees that we can find an $E$-nice basis starting with $\mathcal{G}_{0}^{1}$.
@@ 4.2.3. Case $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
Notice that $\beta_{0}=\beta_{1}$ in this case. Recall Definition 4.1 for $\overline{\Omega}_{1},\overline{\Omega}_{2}$ and $\overline{\Omega}_{j}$. As $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$, we have $[\nu_{C}(\overline{\Omega}_{1})]_{n}=[\nu_{C}(\overline{\Omega}_{2})]_{n}=[\beta_{1}]_{n}$, so that $\{\overline{\Omega}_{1},\overline{\Omega}_{2}\}$ is not an $E$-separated family of level 00. Using Lemma 3.5 and Proposition 3.3 we obtain $\Omega_{1}^{1}=\overline{\Omega}_{1}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0}$ with $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{1})$. This family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is $E$-separated of level 00 but does not satisfy $E$-(b) or $E$-(c) in Definition 4.4, so it is not an $E$-terminal family of level 00. This is why, in this case, we need to start at $l_{0}=1$. By Remark 4.1, we just need to provide an $E$-good family of level $1$.
Consider the $E$-ordered family ${\mathcal{G}}_{1}^{0}=\{\Omega_{1}^{0},\ldots,\Omega_{2{\nu}_{1}}^{0}\}$ given by $\Omega_{j}^{0}=\overline{\Omega}_{j}$ for $1\leq j\leq 2{\nu}_{1}$, which satisfies $\nu(\Omega_{j}^{0})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$. Note that
(23) $[\nu_{C}(\overline{\Omega}_{2j+1})]_{n}=[\nu_{C}(\overline{\Omega}_{2j+2})]_{n}=[(j+1)\beta_{1}]_{n}.$
if $1\leq 2j+1<2j+2\leq 2{\nu}_{1}$. All congruences $[\beta_{1}]_{n},[2\beta_{1}]_{n},\ldots,[\nu_{1}\beta_{1}]_{n}$ are pairwise different, so the unique situation in which two $1$-forms have congruent ${\mathcal{C}}$-orders is described in equation (23). Since $dy\in\tilde{\Delta}_{\beta_{1},<}^{0}$, we have $\Omega_{2j+1}^{0}\in\tilde{\Delta}_{\beta_{1},<}$ by Corollary 3.3. We also have $\Omega_{2j+2}^{0}\in\tilde{\Delta}_{\beta_{1},\leq}$ by Corollary 3.2. We apply Lemma 3.5 and Proposition 3.3 to obtain
@@
for $2j+2\leq 2{\nu}_{1}$. As a consequence, the family ${\mathcal{G}}_{1}^{1}=\{\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1}\}$ is $E$-good of level $1$ (with $E$-(I) being empty). Note that ${\mathcal{L}}_{1}$ is the subset of odd numbers of $\{1,\ldots,2\nu_{1}\}$ whereas ${\mathcal{U}}_{1}$ is its subset of even numbers. Moreover, we obtain $\nu(\Omega_{j}^{1})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$ by construction. As stated above, Remark 4.1 ends our argument.
@@ 4.3. Existence of a nice $E$-basis
Remark 4.3.
The set $\Lambda_{\mathcal{C}}^{E}$ (Definition 2.7) depends on $n,\beta_{0},\beta_{1},\ldots,\beta_{g}$ and indeed it is straightforward to write an algorithm whose input is $(\delta_{x},\delta_{y},n,\beta_{0},\ldots,\beta_{g})$ and whose output is $\Lambda_{\mathcal{C}}^{E}$. The argument is analogous to the one in Remark 3.18.
@@ 5.2. Families of equisingular curves and desingularization
Consider a curve $\Gamma\in\mathcal{C}$, a suitable pair $(\Gamma,E)$ and the setting in Remark 5.3. Let $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ be a a non-constant analytic family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of curves with $\Gamma^{0}=\Gamma$ with suitable parametrizations of the form $\Gamma^{\epsilon}(t)=(t^{\nu({\mathcal{C}})},\sum_{\beta\geq\beta_{0}}a_{\beta}(\epsilon)t^{\beta})$. Define $(\Gamma_{0}^{\epsilon})=(\Gamma^{\epsilon})$ and $(\Gamma_{i}^{\epsilon})$ the strict transform of $(\Gamma^{\epsilon})$ in $\mathcal{X}_{i}$. As $(\Gamma^{\epsilon})$ is non-constant, there exists an integer $\iota(\Gamma^{\epsilon})\geq 1$ such that $\Gamma_{i}^{\epsilon}$ passes through $P_{i}$ for all $0\leq i<\iota(\Gamma^{\epsilon})$, and $\epsilon$ in a neighborhood of 00 but $\Gamma_{\iota}^{\epsilon}$ does not pass through $P_{\iota(\Gamma^{\epsilon})}$ for any $\epsilon$ in a neighborhood of 00.
@@
Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$ for $j\geq 0$ and consider the family $(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ associated to $\iota^{\prime}$ in the proof of Lemma 5.2 with $\Gamma^{0}_{0}=\Gamma$. Since the sliding divisor $\iota(\Gamma_{\iota-1}^{\epsilon})$ is greater than $1$ by definition, we have, by Remark 5.4:
@@ 5.3. Contact orders between curves and vector fields
Given $X\in{\mathfrak{X}}_{E,i}$, with $X(P_{i})=0$ or $i\geq 1$, such that $\Gamma_{i}$ is not $X$-invariant, we consider the flow $(\phi^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of $X$ in a neighborhood of $P_{i}$ and define the curve $\Gamma_{i}^{\epsilon}=\phi^{\epsilon}(\Gamma_{i})$ for $\epsilon\in\mathbb{C}$ in a neighborhood of 00. The pair $(\Gamma_{i}^{\epsilon},E_{i})$ is in the same equisingularity class as $(\Gamma_{i},E_{i})$ for $\epsilon$ in a neighborhood of 00. By blowing-down we obtain the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ induced by $X$, which is an analytic family in the equisingularity class of $\Gamma\in{\mathcal{C}}$.
@@ 6.2.1. Case $E=\emptyset$ or $E=\{x=0\}$
In this case, $l_{0}=0$ for $\mathcal{C}$ and ${E}_{1}=\{x_{1}=0\}$. A terminal family ${\mathcal{T}}_{0}$ of level 00 of $\mathcal{C}$ is $(\Omega_{1},\Omega_{2})=(dx,x^{\delta_{x}}df_{1})$ whereas a terminal family $\tilde{\mathcal{T}}_{0}$ of level 00 of ${\mathcal{C}}_{1}$ is $(\tilde{\Omega}_{1},\tilde{\Omega}_{2})=(dx_{1},x_{1}df_{1,1})$ (see subsection 5.1.1). Of course, $dx,dx_{1}\in\Delta_{n,<}$ and $df_{1},x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, and also

### v3

@@ 4.1. Setting for the case $E\neq\emptyset$
Notice that the only modification is to require that $E$ is invariant, in EE-(i) and EE-(ii).
@@
A family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})\subset\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $l_{0}\leq l\leq g$, is $E$-terminal if
  • EE-(a)
@@
Given an $E$-terminal family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ of level $0\leq l<g$, we define an $E$-initial family ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2\nu_{l+1}}^{0})$ of stage 0 (in level $l+1$) by means of equation (5), exactly as in the case $E=\emptyset$.
@@ 4.2. The method for the case $E\neq\emptyset$
When $E=\emptyset$, the terminal family of level 0 was ${\mathcal{T}}_{0}=(dx,df_{1})$. From this family, applying the results in subsections 3.5.2 and 3.5.3 we ended up with the terminal family ${\mathcal{T}}_{g}$.
@@
Thus, we only need to find an $E$-terminal family of some level $l_{0}$. Actually, it suffices to find an $E$-good family ${\mathcal{G}}_{l_{0}}^{1}$ (with EE-(I) being empty) since in such a case the results in subsection 3.5.3 provide an $E$-terminal family of level $l_{0}$.
@@ 4.2.1. Case $E=\{x=0\}$
which satisfy $\nu(\Omega_{1})=0$, $\nu_{\mathcal{C}}(\Omega_{1})=n$, $\nu(\Omega_{2})=1$, $\nu_{\mathcal{C}}(\Omega_{2})=\beta_{1}+n$. The family $\mathcal{T}_{0}=(\Omega_{1},\Omega_{2})$ is clearly $E$-terminal of level $l_{0}=0$ and thus, we obtain an $E$-nice basis $\mathbf{\Omega}$ by Remark 4.1.
@@ 4.2.2. Case $\delta_{y}=1$ and $\beta_{0}\in(n)$
Notice that $\beta_{0}\in(n)$ implies $\beta_{0}<\beta_{1}$. We have $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$ and hence $\nu_{C}(\overline{\Omega}_{1})\in(n)$, $\nu_{C}(\overline{\Omega}_{2})\in(n)$. We set $l_{0}=0$, $\Omega_{1}^{0}=\overline{\Omega}_{1}$ and $\Omega_{2}^{0}=\overline{\Omega}_{2}$. Applying Lemma 3.5, with $\beta_{l+1}=\beta_{0}$, we obtain $\Omega_{1}^{1}=\Omega_{1}^{0}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{0}),<}^{0}$ where $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{0})$. The family ${\mathcal{G}}_{0}^{1}=(\Omega_{1}^{1},\Omega_{2}^{1})$ is an $E$-good family of level 0. Moreover, it satisfies $\nu(\Omega_{1}^{1})=\delta_{x}$ and $\nu(\Omega_{2}^{1})=1$. Remark 4.1 guarantees that we can find an $E$-nice basis starting with $\mathcal{G}_{0}^{1}$.
@@ 4.2.3. Case $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
Notice that $\beta_{0}=\beta_{1}$ in this case. Recall Definition 4.1 for $\overline{\Omega}_{1},\overline{\Omega}_{2}$ and $\overline{\Omega}_{j}$. As $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$, we have $\nu_{C}(\overline{\Omega}_{1})]_{n}=[\nu_{C}(\overline{\Omega}_{2})]_{n}=[\beta_{1}]_{n}$, so that $\{\overline{\Omega}_{1},\overline{\Omega}_{2}\}$ is not an $E$-separated family of level 0. Using Lemma 3.5 and Proposition 3.3 we obtain $\Omega_{1}^{1}=\overline{\Omega}_{1}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0}$ with $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{1})$. This family ${\mathcal{G}}_{0}^{1}=(\Omega_{1}^{1},\Omega_{2}^{1})$ is $E$-separated of level 0 but does not satisfy EE-(b) or EE-(c) in Definition 4.4, so it is not an $E$-terminal family of level 0. This is why, in this case, we need to start at $l_{0}=1$. By Remark 4.1, we just need to provide an $E$-good family of level $1$.
Consider the $E$-ordered family ${\mathcal{G}}_{1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2{\nu}_{1}}^{0})$ given by $\Omega_{j}^{0}=\overline{\Omega}_{j}$ for $1\leq j\leq 2{\nu}_{1}$, which satisfies $\nu(\Omega_{j}^{0})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$. Note that
(23) $\nu_{C}(\overline{\Omega}_{2j+1})]_{n}=[\nu_{C}(\overline{\Omega}_{2j+2})]_{n}=[(j+1)\beta_{1}]_{n}.$
if $1\leq 2j+1<2j+2\leq 2{\nu}_{1}$. All congruences $\beta_{1}]_{n},[2\beta_{1}]_{n},\ldots,[\nu_{1}\beta_{1}]_{n}$ are pairwise different, so the unique situation in which two $1$-forms have congruent $\mathcal{C}$}-orders is described in equation (23). Since $dy\in\tilde{\Delta}_{\beta_{1},<}^{0}$, we have $\Omega_{2j+1}^{0}\in\tilde{\Delta}_{\beta_{1},<}$ by Corollary 3.3. We also have $\Omega_{2j+2}^{0}\in\tilde{\Delta}_{\beta_{1},\leq}$ by Corollary 3.2. We apply Lemma 3.5 and Proposition 3.3 to obtain
@@
for $2j+2\leq 2{\nu}_{1}$. As a consequence, the family ${\mathcal{G}}_{1}^{1}=(\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1})$ is $E$-good of level $1$ (with EE-(I) being empty). Note that ${\mathcal{L}}_{1}$ is the subset of odd numbers of $\{1,\ldots,2\nu_{1}\}$ whereas ${\mathcal{U}}_{1}$ is its subset of even numbers. Moreover, we obtain $\nu(\Omega_{j}^{1})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$ by construction. As stated above, Remark 4.1 ends our argument.
@@ 4.3. Existence of a nice $E$-basis
4.4. An explicit algorithm
As a byproduct of our work, we present Algorithm 1 in pseudo-code which, given the sequence $(n,\beta_{1},\ldots,\beta_{g})$ of the multiplicity and the Puiseux exponents of an equisingularity class $\mathcal{C}$}, returns the orders $a(1),a(2),\ldots,a(n)$ corresponding to the nice (Apery) basis, $a(i)=\nu_{{\mathcal{C}}}(\Omega_{i})$, for $i=1,\ldots,n$. In particular, the algorithm provides the generic semimodule $\Lambda_{\mathcal{C}}$ and gives Theorem 1.2. Indeed, the powerful properties of our construction allow to describe $\Lambda_{\mathcal{C}}$ without calculating explicitly a nice basis. Notice how, despite the elaborate nature of the previous arguments, the algorithm is straightforward and is coding very simple. Note that the value $b(i)$ is the leading variable of $\Omega_{i}$. The algorithm provides ${\Lambda}_{\mathcal{C}}$ but also gives ${\Lambda}_{\mathcal{C}}^{E}$ for $E=\{x^{\delta_{x}}y^{\delta_{y}}=0\}$ by making the following minor changes:
  • For $\delta_{y}=0$, we consider $a(1)=b(1)=n$, $a(2)=n\delta_{x}+\beta_{1}$ and $b(2)=\beta_{1}$;
  • For $\delta_{y}=1$, we consider $a(1)=n\delta_{x}+\beta_{0}$, $b(1)=\beta_{0}$, $a(2)=n+\beta_{1}$ and $b(2)=\beta_{1}$.
Algorithm 1 Computation of the orders of an Apery basis
Input: $\beta_{0}=n,\beta_{1},\ldots,\beta_{g}$
Values corresponding to $\mathcal{T}_{0}$:
$a(1)\leftarrow n,\,b(1)\leftarrow n,\,a(2)\leftarrow\beta_{1},\,b(2)\leftarrow\beta_{1}$
for $l=0,\ldots,g-1$ do
  Compute the orders corresponding to $\mathcal{G}^{0}_{l+1}$:
  for $k=1,\ldots,n_{l+1}$ do
   for $s=1,\ldots,2\nu_{l}$ do
     $a(2k\nu_{l}+s)\leftarrow\overline{\beta}_{l+1}k+a(s)$
     $b(2k\nu_{l}+s)\leftarrow\beta_{l+1}$
   end for
  end for
  $d\leftarrow\beta_{l+1},\,c\leftarrow 1$
  while $d<\beta_{l+2}\land c\neq 0$ do
   Compute the orders corresponding to $\mathcal{G}^{r+1}_{l+1}$ from those of $\mathcal{G}^{r}_{l+1}$:
   $c\leftarrow 0$
   while $\exists k<s\,\big|(a(s)-a(k)\in(n)\land\max(b(s),b(k))=d)$ do
     $a(s)\leftarrow a(s)+{\mathfrak{n}}(d)-d$
     $b(s)\leftarrow{\mathfrak{n}}(d)$
     if $s\leq n$ then
      $c\leftarrow c+1$
     end if
   end while
   $d\leftarrow\mathfrak{n}(d)$
  end while
end for
Output: $a(1),\ldots,a(n)$, the orders of the nice (Apery) basis
@@ 5.2. Families of equisingular curves and desingularization
Consider a curve $\Gamma\in\mathcal{C}$, a suitable pair $(\Gamma,E)$ and the setting in Remark 5.3. Let $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ be a a non-constant analytic family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of curves with $\Gamma^{0}=\Gamma$ with suitable parametrizations of the form $\Gamma^{\epsilon}(t)=(t^{\nu({\mathcal{C}})},\sum_{\beta\geq\beta_{0}}a_{\beta}(\epsilon)t^{\beta})$. Define $(\Gamma_{0}^{\epsilon})=(\Gamma^{\epsilon})$ and $(\Gamma_{i}^{\epsilon})$ the strict transform of $(\Gamma^{\epsilon})$ in $\mathcal{X}_{i}$. As $(\Gamma^{\epsilon})$ is non-constant, there exists an integer $\iota(\Gamma^{\epsilon})\geq 1$ such that $\Gamma_{i}^{\epsilon}$ passes through $P_{i}$ for all $0\leq i<\iota(\Gamma^{\epsilon})$, and $\epsilon$ in a neighborhood of 0 but $\Gamma_{\iota}^{\epsilon}$ does not pass through $P_{\iota(\Gamma^{\epsilon})}$ for any $\epsilon$ in a neighborhood of 0.
@@
Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$ for $j\geq 0$ and consider the family $(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ associated to $\iota^{\prime}$ in the proof of Lemma 5.2 with $\Gamma^{0}_{0}=\Gamma$. Since the sliding divisor $\iota(\Gamma_{\iota-1}^{\epsilon})$ is greater than 11 by definition, we have, by Remark 5.4:
@@ 5.3. Contact orders between curves and vector fields
Given $X\in{\mathfrak{X}}_{E,i}$, with $X(P_{i})=0$ or $i\geq 1$, such that $\Gamma_{i}$ is not $X$-invariant, we consider the flow $(\phi^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of $X$ in a neighborhood of $P_{i}$ and define the curve $\Gamma_{i}^{\epsilon}=\phi^{\epsilon}(\Gamma_{i})$ for $\epsilon\in\mathbb{C}$ in a neighborhood of 0. The pair $(\Gamma_{i}^{\epsilon},E_{i})$ is in the same equisingularity class as $(\Gamma_{i},E_{i})$ for $\epsilon$ in a neighborhood of 0. By blowing-down we obtain the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ induced by $X$, which is an analytic family in the equisingularity class of $\Gamma\in{\mathcal{C}}$.
@@ 6.2.1. Case $E=\emptyset$ or $E=\{x=0\}$
In this case, $l_{0}=0$ for $\mathcal{C}$ and ${E}_{1}=\{x_{1}=0\}$. A terminal family ${\mathcal{T}}_{0}$ of level 0 of $\mathcal{C}$ is $(\Omega_{1},\Omega_{2})=(dx,x^{\delta_{x}}df_{1})$ whereas a terminal family $\tilde{\mathcal{T}}_{0}$ of level 0 of ${\mathcal{C}}_{1}$ is $(\tilde{\Omega}_{1},\tilde{\Omega}_{2})=(dx_{1},x_{1}df_{1,1})$ (see subsection 5.1.1). Of course, $dx,dx_{1}\in\Delta_{n,<}$ and $df_{1},x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, and also